find two consecutive numbers such that five-eights of the smaller number equals three-fifths of the larger number.
Answer by mananth(16949)

let the numbers be x & x+1
...
5x/8 = 3(x+1)/5
25x=24(x+1)
25x= 24x+24
25x-24x=24
x= 24
..
The numbers are 24 & 25
...
C H E C K
5*24/8 =3(24+1)/5
15 = 15
....
m.ananth@hotmail.ca